Order: 250 mL packed red blood cells - run over 3 hrs.

Calculate the volume per hour.

Calculate the volume per minute using standard blood tubing
Answer #1

Prescribed dose =250 ml for 3 hours

Volume per hour = total volume /total time

=250/3= 83.3 ml/ hour

Volume per minute= volume per hour/60

83.3/60= 1.39 ml/ minute